Meeta can row her boat at 5km per hour in sill water.If she takes 1hr more to row the boat 5.25km upstream than to return downstream,find the speed of the stream.
Answer by stanbon(75887)

Meeta can row her boat at 5km per hour in still water. If she takes 1hr more to row the boat 5.25km upstream than to return downstream, find the speed of the stream.
--------
Upstream DATA:
rate = 5 - s ; distance = 5.25 km ; time = d/r = 5.25/(5-s) hrs.
--------------------
Downstream DATA:
rate = 5+s ; distance = 5.25 km ; time = d/r = 5.25/(5+s) hrs
--------------------
Equation:
time - time = 1 hr
5.25/(5-s) - 5.25/(5+s) = 1
-----
5.25*5 + 5.25s - 5.25*5 + 5.25s = 25-s^2
--------
10.5s = 25-s^2
--------
s^2 + 10.5s - 25 = 0
-----
Factor:
(s-2)(s+12.5) = 0
------
Positive solution:
s = 2 km/hr (speed of the stream)
=====================
Cheers,
Stan H.
=====================